OpenStudy (anonymous):

-9/10 - (-9/10)= ? this lesson is adding and subtracting with like denominators, but what do u do with the fraction in ( ) ?

OpenStudy (anonymous):

=0.

OpenStudy (anonymous):

sign of second fraction changes to + result is 0

OpenStudy (anonymous):

Look at it like this. -9/10 -1(-9/10) So -1*-9/10 is 9/10 -9/10+9/10=0

OpenStudy (anonymous):

thank u:)

OpenStudy (anonymous):

Welcome man, just remember when you have a - in front of parentheses like this... -(abunchofnumbershere) its like -1*abunchofnumbershere=-abunchofnumbershere

OpenStudy (anonymous):

how bout on like this , 25/14 + (-13\14)=

OpenStudy (anonymous):

When you multiply a positive by a negative you get a negative right? Or when dealing with the + in front of the (-13/14) to alleviate confusion they have to add in the parentheses or it would look like this... 25/14+-13/14...that can be confusing. Think of the parentheses as a shields up, protecting the contents of the parentheses from that pesky + sign.

OpenStudy (anonymous):

well the problem is actually -25/14 +(-13\14)= ......the answer i got was -2 5/7 simplied , but not sure if that is correct?

OpenStudy (anonymous):

The answer is -19/7 7 will go into 19 twice with 5 remaining -2 5/7 is right

OpenStudy (anonymous):

ok thanks alot